At a furniture store, a furniture was sold for P475 less than a slide table and P278 more than a kitchen dish cabinet. The total cost of the 3 items was P4175. What was the price of the slide table?
Answer by addingup(3677)

furniture: f
Slide table: f+475
kit cabinet: f-278
f+f+475+f-278 = 4175
3f+197 = 4175
3f = 3978
f = 1326
slide table: 1326+475 = 1801
kit cabinet: 1326-278 = 1048
-----------------------------
Check:
1326+1801+1048 = 4175 Correct